# Multi-agent Adaptive Mechanism Design

### v1

@@ 4. Adaptive Mechanism Design
In this section, we study the problem of adaptive mechanism design where the principal has no knowledge of the nature or the players. Starting from oblivion, the principal’s aim is to dynamically design mechanisms that eventually converge to the optimum, while ensuring players’ truthfulness along the convergence process. At the same time, the principal would attempt to converge as fast as possible, so as to minimize the regret. Define the principal’s (empirical cumulative) regret after $T$ rounds as $\sum_{t=1}^{T}\sum_{i=1}^{N}(R_{it}-R^{*}_{it})$, where $R^{*}_{it}$ is the optimal payment ($\mathbb{E}[R^{*}_{it}]=c$ as shown by Theorem 3.2).
We would demonstrate that the bottleneck of the adaptive mechanism design problem is the difficulty to learn the conditional distribution $p^{*}(x_{j}\mid x_{i})$. Surrounding this observation, we design a statistically optimal adaptive mechanism that achieves regret guarantee with matching lower bounds. We present our algorithm, ”Distributionally Robust Adaptive Mechanism” (DRAM), in Algorithm 1. The algorithm maintains truthfulness and reduces cost by designing a sequence of distributionally robust mechanisms with shrinking ambiguity parameter $\eta$. The ambiguity parameter tracks the accuracy of the principal’s estimation of nature and players’ behavior at each round.
In DRAM, the entire $T$ tasks are divided into two phases: warm-start phase and adaptive phase. As suggested by Theorem 3.7 (see discussions after Theorem 3.7), initially, when the ambiguity level is above a certain threshold $\tilde{\eta}$, no mechanism can ensure truthfulness for all distributions in the ambiguity set. The warm-start phase is designed to reduce ambiguity below that threshold. In this phase, the principal would use the true label $Y_{t}$ for verification. This phase lasts $O(\log\log T)$ tasks, so cost is controlled even when true label might be expensive to get. Then, the principal moves to the adaptive phase. As the principal obtains more data, it constructs a more accurate estimation. This allows it to design a mechanism with decreasing $\eta$ and therefore reduce cost of robustness. The ambiguity parameter $\eta$ shrinks at a proper rate, so as to make sure truthfulness is preserved with high probability for each round.
During the entire algorithm, the principal uses players’ past report to estimate $p^{*}(x_{j}\mid x_{i})$, which is the posterior distribution of reference player’s observation $x_{j}$ conditional on the focal player $x_{i}$. Note that the principal only has players’ reports $z$ but not the true observation $x$. This means truthfulness must be guaranteed at all times to ensure the estimation fidelity. This adds another evidence on the necessity of truthfulness - it is not only necessary for optimal downstream aggregation tasks (Theorem 2.1), but also necessary for cutting costs.
Input : ambiguity threshold $\tilde{\eta}$; failure tolerance $\varepsilon$; lower bound on observation frequency $0<\rho<\min_{i,x\in\mathcal{Y}}\mathbb{P}(X_{i}=x)$.
@@
For each player $i$, assign a corresponding reference player $j$.
@@
   Deploy the fact-checking mechanism $R_{i}=\mathbf{1}\{Z_{it}=Y_{t}\}$ for each player $i$.
@@
  For each player $i$, set their safety margin
@@
  Compute the mechanism $R_{ik}$ for each player by solving Eq.(4) with parameter $\hat{p}_{it}$ and $\delta_{ik}$.
@@
Inputs. Out of the three input variables, the failure tolerance is one that can be decided arbitrarily, and the other two depend on the players we work with. The ambiguity threshold for all players is defined as $\tilde{\eta}=\min_{i}\tilde{\eta_{i}}$, where the player-specific threshold $\tilde{\eta}_{i}$ is computed according to Theorem 3.7:
@@
Warm-start phase. The main objective of the warm-start phase is to reduce the principal’s ambiguity below the threshold suggested by Theorem 3.7 so that distributionally robust mechanisms can be applied. There are multiple approaches to reduce ambiguity, and here the principal learns $p^{*}(x_{j}\mid x_{i})$ by collecting truthful reports from players. The main challenge is to ensure truthfulness when peer prediction mechanisms do not work. We achieve this by using true label verification. Suppose the principal can now obtain the correct label $Y_{t}$ from an external expert. This expert verification might require a cost much higher than the total observation cost per round $Nc$, therefore it is costly to apply it for too many rounds. The principal can now compare players’ reports with ground truth and reward according to a fact-checking mechanism $R_{it}(Z_{it},Y_{t})$.
@@
Recall the assumption that each true label $y$ appears with uniformly bounded probability $\underline{p}\leq p_{Y}(y)\leq\overline{p}$. If for all $y\in\mathcal{Y}$ and $x\neq y$, player $i$’s skill $p_{i}$ satisfies the diagonal dominance property:
@@
then the simple fact-checking rule $R_{it}(Z_{it},Y_{t})=\mathbf{1}\{Z_{it}=Y_{t}\}$ guarantees player $i$’s truthfulness.
The diagonal dominance condition essentially assumes that players are more likely to obtain the correct observation than to make a mistake. Therefore, any lying behavior would decrease the probability of the report being correct; thus, the player should be incentivized to tell the truth. We note that it is generally impossible to design a fact-checking rule that guarantees truthfulness for arbitrary skill distribution $p_{i}$ and $p_{Y}$. It is shown in (lambert2011elicitation) that if a player’s observation has overlaps, i.e., the player can have the same observation under two different labels, then there always exists an adversarial prior under which a fact-checking mechanism fails.
Adaptive phase. After the ambiguity is lower than the threshold, the principal moves into the adaptive phase. This phase divides the entire time horizon into epochs, with each epoch double the size of the previous one. In total, we would have $O(\log T)$ epochs.
At the beginning of each epoch, the principal would call two oracles: i) an offline estimation oracle for the reference distribution $p^{*}(x_{j}\mid x_{i})$ (in Algorithm 1 it is the empirical distribution estimator), and ii) an optimization oracle that computes the distributionally robust mechanisms by solving Eq.(4). Then, the principal would use the same produced mechanism throughout the entire epoch, and no further computation is needed. This indicates DRAM is also computationally efficient with $O(N\log T)$ total calls to both oracles.
@@
Consider the sequential mechanism design problem with $N$ players and $T$ rounds. With probability at least $1-\varepsilon$, Algorithm 1 simultaneously achieves:
  • truthfulness is guaranteed for all $N$ players in all $T$ rounds.
@@
Theorem 4.2 recovers the $O(\sqrt{T})$ terms typically seen in bandits and online learning literature (lattimore2020bandit). In fact, oracle calls can be further reduced to $O(N\log\log T)$ when $T$ is known. In DRAM, we use the classical doubling trick (cesa2006prediction) from the online learning literature. This trick does not require exact knowledge of the number of tasks. (Although we need to know the magnitude $\log(T)$ to compute epochwise ambiguity parameter $\eta_{k}$s.) The epoch schedule $\tau_{k}-\tau_{k-1}=T^{1-2^{-(k-1)}}\tau$ (similar to (cesa2014regret)) maintains the same regret guarantee, while requiring even fewer oracle calls $O(\log\log T)$.
@@
Consider replacing the epoch schedule in Algorithm 1 with $\tau_{k}-\tau_{k-1}=T^{1-2^{-(k-1)}}\tau$. With probability at least $1-\varepsilon$, the principal simultaneously guarantees truthfulness across all rounds for all players, and the expected total regret maintains the same upper bound
@@ 4.1. Regret Lower Bound of Adaptive Mechanism Design
We now show that DRAM is statistically optimal up to logarithmic factors. In particular, we prove a matching lower bound demonstrating that any policy which guarantees truthfulness with high probability must incur regret of order at least $\Omega(N\sqrt{T})$. Throughout this section, we consider the general setting described in Remark 1, where players’ types (observations) are directly sampled from an unknown joint distribution $p_{\boldsymbol{X}}$ over the type space $\mathcal{Y}^{N}$.
@@
Consider the sequential mechanism design problem with $N$ players and $T$ rounds. Fix any failure tolerance $\varepsilon\in(0,1/4)$. For any (possibly randomized) non-anticipating reward policy that guarantees truthfulness across all players and rounds with probability at least $1-\varepsilon$, there exists a type distribution $p_{\boldsymbol{X}}\in\Delta(\mathcal{Y^{N}})$ under which, with probability at least $1-\varepsilon$, the total regret is at least
@@
The result reveals that the inherent difficulty of adaptive mechanism design lies in learning players’ conditional beliefs, namely the posterior distributions $p^{*}(x_{j}\mid x_{i})$ that govern incentives. Because the lower bound is derived via a two-point argument, it does not explicitly depend on the alphabet size $d=|\mathcal{Y}|$. It remains to be studied the optimality of . However, since estimating a discrete distribution over $\mathcal{Y}$ incurs a minimax risk of order $\sqrt{d/T}$ (han2015minimax), we conjecture that the regret bound achieved by DRAM is also optimal in its dependence on $d$, up to logarithmic factors.
@@ 4.2. Extension to General Estimator
An important observation of DRAM is that the estimation oracle and the optimization oracle are separate modules. They are connected via the ambiguity parameter $\eta_{k}$, which measures the distance between the estimated distribution and the actual one. This means that DRAM is flexible with estimators, so long as its estimation could satisfy the requirement in Eq.(5). Therefore, the empirical estimator could be swapped with any other distribution estimator that may better exploit and reflect the underlying structures of players’ skills. Based on this, we propose the algorithm $\texttt{DRAM}+$, which works with general distribution estimators.
@@
This estimation guarantees $\eta_{\varepsilon}(t)$ should monotonically decrease with $t$, and monotonically increase with failure probability $\varepsilon$. It bounds the gap between $\hat{q}_{t}$ and $q$ under a dataset size of $t$. Such a bound is commonly seen in the probably approximately correct (PAC) framework of statistical learning, where a good estimator should achieve a lower gap with higher probability and lower $t$.
Now we introduce $\texttt{DRAM}+$ (Algorithm 2), which modifies DRAM to work with general discrete distribution estimators in Definition 4.5. Compared to DRAM, the main difference of $\texttt{DRAM}+$ lies in the epoch schedule and ambiguity parameters. In $\texttt{DRAM}+$, we don’t restrict how epoch schedules are designed. This may lead to suboptimal schedules. Generally, one should aim for a geometric epoch schedule, as this typically results in the best possible bounds and only $m=O(\log T)$ epochs. The impact of schedule is studied in (perchet2016batched; besson2018doubling), from the perspective of batched feedback and doubling trick, respectively. Moreover, the ambiguity parameters now follow the guarantee $\eta_{\varepsilon}(t)$, in order to ensure truthfulness holds with high probability.
@@
For each player $i$, assign a corresponding reference player $j$.
@@
Consider the sequential mechanism design problem with $N$ players and $T$ rounds. With probability at least $1-\varepsilon-N(d+1)\cdot\sum_{k=1}^{m}\exp(-\rho\tau_{k-1}/8)$, Algorithm 2 simultaneously achieves:
  • truthfulness is guaranteed for all $N$ players in all $T$ rounds.
@@
Compared to Theorem 4.6, an additional overhead $N(d+1)\cdot\sum_{k=1}^{m}\exp(-\rho\tau_{k-1}/8)$ appears in the high probability guarantee in the Theorem. This is due to a failed event when a player does not observe a certain label $x_{i}$ for enough number of times, and there are not enough data to recover $p^{*}(\cdot|x_{i})$. This failed case is inherent and universal for whatever estimator the principal uses. Without a closed-form description of the estimation gap, we cannot absorb this probability into $\varepsilon$, like what we did in Theorem 4.6. (See Appendix A for more details.) Nonetheless, with appropriately chosen schedule (such as $\tau_{k}-\tau_{k-1}=2^{k-1}\tau$), this exponential term should be dominated by $\varepsilon$.
The central interpretation of $\texttt{DRAM}+$ and Theorem 4.6 is that any estimation guarantees for discrete distribution can be immediately translated to mechanism regret guarantees. When the principal starts with assumptions or prior beliefs on the distribution $p^{*}(x_{j}\mid x_{i})$, it can apply a different estimator to exploit that knowledge.
We note that the design of DRAM and $\texttt{DRAM}+$ is conceptually related to the multi-armed bandit algorithm from (simchi2022bypassing). In fact, both DRAM and $\texttt{DRAM}+$ are partly inspired by their insight that an online contextual bandit problem can be reduced to an offline regression problem. Compared to the classical multi-armed bandit, the adaptive mechanism design problem is, in some sense, both simpler and harder. On the one hand, the core challenge in bandit problems lies in the exploration–exploitation trade-off, since the arm-sampling policy in previous rounds affects observed data distributions. From that perspective, the mechanism design problem is simpler than bandits, since the underlying distributions remain unaffected by the principal’s mechanism decision as long as players behave truthfully. On the other hand, when participants are rational, incentivizing truthfulness and doing it optimality is nontrivial. Players’ incentives and skills are unknown, and any deviation can cause unpredictable dynamics. In contrast, there are no incentives involved in bandits, and each arm always gives truthful feedback. Nonetheless, the connection to (simchi2022bypassing), together with Theorem 4.6, suggests that a principled reduction from online mechanism design to offline learning may indeed be possible.

### v2

@@ 4. Adaptive Mechanism Design
In this section, we study the problem of adaptive mechanism design where the principal has no initial knowledge. Define the principal’s (empirical cumulative) regret after $T$ rounds as $\sum_{t=1}^{T}\sum_{i=1}^{N}(R_{it}-R^{*}_{it})$, where $R^{*}_{it}$ is the optimal payment ($\mathbb{E}[R^{*}_{it}]=c$ as shown by Theorem 3.2). Starting from oblivion, the principal’s aim is to minimize regret, while ensuring agents’ truthfulness.
We present our algorithm, “Distributionally Robust Adaptive Mechanism” (DRAM), in Algorithm 1. The algorithm maintains truthfulness and reduces cost by designing a sequence of distributionally robust mechanisms with shrinking ambiguity parameter $\eta$. The ambiguity parameter tracks the accuracy of the principal’s estimation at each round.
Input : ambiguity threshold $\tilde{\eta}$; failure tolerance $\varepsilon$; lower bound on observation frequency $0<\rho<\min_{i,x\in\mathcal{Y}}\mathbb{P}(X_{i}=x)$; upper bound on observations $\max_{y}\mathbb{P}(X_{i}=y)<\gamma_{(i)}<1$.
@@
For each agent $i$, assign a corresponding reference agent $j$.
@@
   Deploy the fact-checking mechanism $R_{i}=\mathbf{1}\{Z_{it}=Y_{t}\}$ for each agent $i$.
@@
  For each agent $i$, set their safety margin
@@
  Compute the mechanism $R_{ik}$ for each agent by solving Eq.(3) with parameter $\hat{p}_{it}$ and $\delta_{ik}$.
@@
In DRAM, the entire horizon is divided into two phases: warm-start phase and adaptive phase. As suggested by Theorem 3.7, our scheme fails when the ambiguity level is above a certain threshold $\tilde{\eta}$. Therefore, the warm-start phase is designed to reduce ambiguity below that threshold. In this phase, the principal would use ground truth $Y_{t}$ for verification. Then, the principal moves to the adaptive phase, which is split into epochs. At the beginning of each epoch, the principal uses the empirical distribution for estimation. As the principal obtains more data, estimation is more accurate. This allows it to design a mechanism with decreasing $\eta$ and therefore reduce additional cost for robustness. The ambiguity parameter $\eta$ shrinks at a proper rate, so as to make sure truthfulness is preserved with high probability for each round.
During the entire algorithm, the principal uses agents’ past report to estimate $p^{*}(x_{j}\mid x_{i})$, which is the posterior distribution of reference agent’s observation $x_{j}$ conditional on the focal agent $x_{i}$. Note that the principal only has agents’ reports $z$ but not the true observation $x$. This means truthfulness must be guaranteed at all times for estimation fidelity. This adds another evidence on the necessity of truthfulness in addition to Theorem 2.1.
Inputs. Out of all the input variables, failure tolerance can be decided arbitrarily, and the rest depend on the agents. The ambiguity threshold for all players is defined as $\tilde{\eta}=\min_{i}\tilde{\eta_{i}}$, where the agent-specific threshold $\tilde{\eta}_{i}$ is computed according to Theorem 3.7:
@@
Warm-start phase. The main objective of the warm-start phase is to reduce the principal’s ambiguity below the threshold suggested by Theorem 3.7, so that distributionally robust mechanisms can be applied. There are multiple approaches to reduce ambiguity, and here the principal learns $p^{*}(x_{j}\mid x_{i})$ by collecting truthful reports from agents. We incentivize truthfulness in this phase by using ground truth verification. Suppose the principal can now obtain the ground truth $Y_{t}$ from an external expert. With ground truth available, The principal can compare reports with it, and then reward according to a fact-checking mechanism $R_{it}(Z_{it},Y_{t})$. This phase lasts $O(\log\log T)$ tasks, so cost is controlled even when ground truth is expensive.
@@
Recall the assumption that each true label $y$ appears with uniformly bounded probability $\underline{p}\leq p_{Y}(y)\leq\overline{p}$. If for all $y\in\mathcal{Y}$ and $x\neq y$, agent $i$’s skill $p_{i}$ satisfies the diagonal dominance property:
@@
then the simple fact-checking rule $R_{it}(Z_{it},Y_{t})=\mathbf{1}\{Z_{it}=Y_{t}\}$ guarantees agent $i$’s truthfulness.
The diagonal dominance condition essentially assumes that agents are more likely to obtain the correct observation than to make a mistake. Therefore, any lying behavior would decrease the probability of the report being correct; thus, the agent should be incentivized to tell the truth. We note that it is generally impossible to design a fact-checking rule that guarantees truthfulness for arbitrary skill distribution $p_{i}$ and $p_{Y}$. It is shown in (Lambert, 2011) that if a agent’s observation has overlaps, i.e., the agent can have the same observation under two different labels, then there always exists an adversarial prior under which a fact-checking mechanism fails.
Adaptive phase. After the ambiguity is lower than the threshold, the principal moves onto the adaptive phase. This phase divides the entire time horizon into epochs, with each epoch double the size of the previous one. In total, we would have $O(\log T)$ epochs.
At the beginning of each epoch, the principal would call two oracles: i) an offline estimation oracle for the reference distribution $p^{*}(x_{j}\mid x_{i})$ (in Algorithm 1 it is the empirical distribution estimator), and ii) an optimization oracle that computes the distributionally robust mechanisms by solving Eq.(3). Then, the principal would use the same produced mechanism throughout the entire epoch, and no further computation is needed. This indicates DRAM is computationally efficient with $O(N\log T)$ total calls to both oracles.
@@
Consider the sequential mechanism design problem with $N$ agents and $T$ rounds. With probability at least $1-\varepsilon$, Algorithm 1 simultaneously achieves:
  • truthfulness is guaranteed for all $N$ agents in all $T$ rounds.
@@
Theorem 4.2 recovers the $O(\sqrt{T})$ terms typically seen in bandits and online learning literature (Lattimore and Szepesvári, 2020). In fact, oracle calls can be further reduced to $O(N\log\log T)$ when $T$ is known. In DRAM, we use the classical doubling trick (Cesa-Bianchi and Lugosi, 2006) from the online learning literature. This trick does not require exact knowledge of the number of tasks. (Although we need to know the magnitude $\log(T)$ to compute epochwise ambiguity parameter $\eta_{k}$s.) The epoch schedule $\tau_{k}-\tau_{k-1}=T^{1-2^{-(k-1)}}\tau$ (similar to (Cesa-Bianchi et al., 2014)) maintains the same regret guarantee, while requiring even fewer oracle calls $O(\log\log T)$.
@@
Consider replacing the epoch schedule in Algorithm 1 with $\tau_{k}-\tau_{k-1}=T^{1-2^{-(k-1)}}\tau$. With probability at least $1-\varepsilon$, the principal simultaneously guarantees truthfulness across all rounds for all agents, and the expected total regret maintains the same upper bound
@@ 4.1. Regret Lower Bound of Adaptive Mechanism Design
We now show that DRAM is statistically optimal up to logarithmic factors. In particular, we prove a matching lower bound demonstrating that any policy which guarantees truthfulness with high probability must incur regret of order at least $\Omega(N\sqrt{T})$.
@@
Consider the sequential mechanism design problem with $N$ agents and $T$ rounds. Fix any failure tolerance $\varepsilon\in(0,1/4)$. For any (possibly randomized) non-anticipating reward policy that guarantees truthfulness across all agents and rounds with probability at least $1-\varepsilon$, there exists a type distribution $p_{\boldsymbol{X}}\in\Delta(\mathcal{Y^{N}})$ under which, with probability at least $1-\varepsilon$, the total regret is at least
@@
The results together with its proof (see Appendix A) reveals that the regret bottleneck of adaptive mechanism design is the difficulty in learning players’ conditional beliefs, namely the posterior distributions $p^{*}(x_{j}\mid x_{i})$ that govern incentives. Because the lower bound is derived via a two-point argument, it does not explicitly depend on the alphabet size $d=|\mathcal{Y}|$. However, since estimating a discrete distribution over $\mathcal{Y}$ incurs a minimax risk of order $\sqrt{d/T}$ (Han et al., 2015), we conjecture that the regret bound achieved by DRAM is also optimal in its dependence on $d$, up to logarithmic factors.
@@ 4.2. Extension to General Estimator
An important observation of DRAM is that the estimation oracle and the optimization oracle are decoupled. They are connected via the ambiguity parameter $\eta_{k}$, which measures the distance between the estimated distribution and the actual one. This means that DRAM is flexible with estimators, so long as its estimation could satisfy the requirement in Eq.(4). Therefore, the empirical estimator could be swapped with any other distribution estimator that may better exploit and reflect the underlying structures of agents’ skills. Based on this, we propose the algorithm $\texttt{DRAM}+$ working with general distribution estimators.
@@
This estimation guarantees $\eta_{\varepsilon}(t)$ should monotonically decrease with $t$, and monotonically increase with failure probability $\varepsilon$. Such a bound is commonly seen in the probably approximately correct (PAC) framework of statistical learning, where a good estimator should achieve a lower gap with higher probability and lower $t$.
Now we introduce $\texttt{DRAM}+$ (Algorithm 2), which modifies DRAM to work with general discrete distribution estimators in Definition 4.5. In $\texttt{DRAM}+$, we don’t restrict how epoch schedules are designed. Generally, one should aim for a geometric epoch schedule, as this typically results in the best possible bounds and only $m=O(\log T)$ epochs. Moreover, the ambiguity parameters now follow the guarantee $\eta_{\varepsilon}(t)$, in order to ensure truthfulness holds with high probability.
@@
For each agent $i$, assign a corresponding reference agent $j$.
@@
Consider the sequential mechanism design problem with $N$ agents and $T$ rounds. With probability at least $1-\varepsilon-N(d+1)\cdot\sum_{k=1}^{m}\exp(-\rho\tau_{k-1}/8)$, Algorithm 2 simultaneously achieves:
  • truthfulness is guaranteed for all $N$ agents in all $T$ rounds.
@@
Compared to Theorem 4.6, an additional overhead $N(d+1)\cdot\sum_{k=1}^{m}\exp(-\rho\tau_{k-1}/8)$ appears in the high probability guarantee. This is due to a failed event when a player does not observe a certain label $x_{i}$ for enough number of times, and there are not enough data to recover $p^{*}(\cdot|x_{i})$. This failed event is universal, and without a closed-form description of the estimation gap, we cannot merge this probability with $\varepsilon$. Nonetheless this term decays exponentially, and, with appropriately chosen schedule (such as $\tau_{k}-\tau_{k-1}=2^{k-1}\tau$), it should be dominated by $\varepsilon$.
The central interpretation of $\texttt{DRAM}+$ and Theorem 4.6 is that any estimation guarantees for discrete distribution can be immediately translated to mechanism regret guarantees. This suggests that a principled reduction from online mechanism design to offline learning may indeed be possible, similar to the reduction from contextual bandits to offline estimation in (Simchi-Levi and Xu, 2022).
Compared to the classical multi-armed bandit, the adaptive mechanism design problem is, in some sense, both simpler and harder. On the one hand, the core challenge in bandit problems lies in the exploration–exploitation trade-off, since the arm-sampling policy in previous rounds affects observed data distributions. From that perspective, the mechanism design problem is simpler than bandits, since the underlying distributions remain unaffected by the principal’s mechanism decision as long as agents behave truthfully. On the other hand, when participants are rational, incentivizing truthfulness in an optimal way is nontrivial. Agents’ incentives and skills are unknown, and any deviation can cause unpredictable dynamics. In contrast, there are no incentives involved in bandits, and each arm always gives truthful feedback.